Please help me in answering this problem:
The unit measure for atmospheric pressure is KPa (Kilo-Pascals). At sea level the standard atmospheric pressure is 101.3 KPa. For every 1000 feet increase in altitude atmospheric pressure decreases by 1.8 KPa.
What is the atmospheric pressure at 50,000 feet?
I have tried to answer the question, my understanding to the question is what is the value of t 50,000 and the common difference d is 1000 but I didn't know if I should consider the first term (a)= 101.3 or 101.3 - 1.8 = 99.5 KPa
Thank you

Answer by richwmiller(17219)

At sea level, the height is 0 and the pressure is 101.3
At 1000 ft, the height is 1000 and the pressure is 99.5
The first time we lose 1.8 is at 1000 ft or 1
(0,101.3)
(1, 99.5)
(2, 97.7)
So we lose 1.8 for each and every 1
The common difference is -1.8 for every 1000 ft which we will consider 1
The first time we lose 1.8 is at 1000 ft or 1
we lose 1.8 for every 1000 ft
If you work out the slope and y intercept
y = 101.3-1.8x
y=101.3-1.8*50
y=11.3 KPa at 50000 ft
Your problem stems from starting with 0 at sea level and not 1000
The linear equation is still y = 101.3-1.8x if we use the points (1, 99.5) and
(2, 97.7) where x=50 and y=11.3
An = a1 + (n - 1)d
A50 = 99.5 -1.8(49)
We still get 11.3 at 50000
A50=11.3 KPa
so A1 = 99.5 and the difference is -1.8